# Motion Consistency Model: Accelerating Video Diffusion with Disentangled Motion-Appearance Distillation

### v1

@@ 3.2 Disentangled motion consistency distillation
We examine all choices in Sec. 4.3, and find that the learnable representation works the best.
@@ 4 Experiments
Figure 6: MCM frame quality improvement results using different image datasets with ModelScopeT2V [62] teacher. Our MCM effectively adapts to different distributions with 4 steps. Corresponding video results are in supplementary materials.
@@ Appendix A Related work
Video generation. Early methods mainly rely on generative adversarial networks (GANs) [58, 6, 67, 53], which struggled with scalability and generalization to out-of-distribution domains. Building on the recent success of image diffusion models [22, 45], several methods leverage diffusion models for video generation [21, 24, 9, 62, 19, 65]. This paper focuses on two popular text-to-video diffusion models: ModelScopeT2V [62] and AnimateDiff [19]. ModelScopeT2V [62] inserts spatio-temporal blocks into 2D backbones and trains the whole backbone end-to-end, ensuring adaptability to varying frame lengths with smooth motions. AnimateDiff [19] introduces a plug-and-play motion module that learns motion priors from real-world videos and can be applied to personalized image diffusion models for video generation.

### v2

@@ 3.2 Disentangled motion consistency distillation
We examine all choices in Sec. 4.3, and find that the learnable representation works the best, enabling effective motion consistency modeling while preserving frame appearance quality.
@@ 4 Experiments
Figure 6: MCM frame quality improvement results using different image datasets with ModelScopeT2V [62] teacher. The first and last frames are sampled for visualization. Our MCM effectively adapts to different distributions with 4 steps. Corresponding video results are in supplementary materials.
@@ 4.1 Comparisons with distilled video diffusion models
Figure 7: Our MCM allows for high-resolution video generations with different aspect ratios, using 4 sampling steps. The left three videos (576x1024) are in portrait format, the top right two videos (768x512) are in landscape format, and the bottom right two videos (768x768) are in square format.
Figure 8: Our MCM can incorporate ControlNet [75] to enable pose-conditioned video generation. The videos are generated using 4 sampling steps.
@@ 4.3 Ablation study
4.4 Applications
We demonstrate that our proposed MCM is versatile and applicable to various tasks, including high-resolution video generation with different aspect ratios and conditioned video generation. In Fig. 7, MCM generates high-resolution videos up to 576x1024 pixels and supports various aspect ratios such as square, portrait, and landscape formats. Moreover, by incorporating controllable modules, MCM enables conditional video generation. For example, in Fig. 8, integrating ControlNet [75] allows for pose-conditioned video generation. Notably, both videos are generated using only 4 sampling steps, highlighting the effectiveness and versatility of our MCM.
@@ Appendix A Related work
Video generation. Early methods mainly rely on generative adversarial networks (GANs) [58, 6, 68, 53], which struggled with scalability and generalization to out-of-distribution domains. Building on the recent success of image diffusion models [22, 45], several methods leverage diffusion models for video generation [21, 24, 9, 62, 19, 65, 66, 74]. This paper focuses on two popular text-to-video diffusion models: ModelScopeT2V [62] and AnimateDiff [19]. ModelScopeT2V [62] inserts spatio-temporal blocks into 2D backbones and trains the whole backbone end-to-end, ensuring adaptability to varying frame lengths with smooth motions. AnimateDiff [19] introduces a plug-and-play motion module that learns motion priors from real-world videos and can be applied to personalized image diffusion models for video generation.